OpenStudy (anonymous):

If 2 x^2+ 2 x + xy = 3 and y( 3 ) = -7, find y'( 3 ) by implicit differentiation.

OpenStudy (anonymous):

help anyone???

OpenStudy (anonymous):

y = (3 - 2x - 2x^2)/x Then differentiate? I don't remember the terminology "implicit differentiation"

OpenStudy (anonymous):

its just like finding the derivative

OpenStudy (anonymous):

but just looking at the y term as a function f(x)

OpenStudy (anonymous):

i think....

OpenStudy (anonymous):

Using implicit differentiation, I suspect this has the form:\[y \prime = - \frac{2 + 4x + y}{x}\]Use the point you have to find the value for y and solve for y'.

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Think about it this way, let y be a function of x. Then we have f(x, y(x)). When we take the derivative of it, we take D{f(x,y)}, so we get: 4x + 2 + D{xy} = D{3} -> D{xy} = -4x - 2 Then, remember the product rule: D{xy} = y + xy', so we are left with: y + xy' = -4x - 2 -> y' = -(4x + 2 + y)/x

OpenStudy (anonymous):

ok i see....so i use y=3 and input this into y prime?

OpenStudy (anonymous):

Yeah, you have the point (3,-7). Put x = 3 and y = -7 and solve for y'

OpenStudy (anonymous):

how does y( 3 ) = -7 come into play?

OpenStudy (anonymous):

-7/3...thanks

OpenStudy (anonymous):

:-) No problem.

OpenStudy (anonymous):

\[2 x^2+ 2 x + xy = 3\implies 4x+2+y+xy'=0 \]\[\implies y'=-\frac{4x+2+y}{x}~since~y(3)=-7 \implies y'(3)=-\frac{4(3)+2+(-7)}{3}=\frac{-7}{3}\]

OpenStudy (anonymous):

Got the request a little late, hope my input is worthwhile.